What is the rule for the composite transformation formed by a translation 2 units to the left and 3 units up, followed by a 90 degree rotation counterclockwise?
Found 2 solutions by Edwin McCravy, stanbon:
Answer by Edwin McCravy(20086)

What is the rule for the composite transformation formed by a translation 2 units to the left and 3 units up, followed by a 90 degree rotation counterclockwise?
To translate 2 units to the left, replace x by (x+2) in the right side.

To then translate 3 units up, add +3 to the right side. 

To do a 90 degree rotation counterclockwise, replace x by y, 
and y by -x, and solve for y.

Edwin

Answer by stanbon(75887)

What is the rule for the composite transformation formed by a translation 2 units to the left and 3 units up, followed by a 90 degree rotation counterclockwise?
========
2 to the left: x goes to x+2
3 up: y goes to y+3
90 degree rotation counterclockwise: (x,y) goes to (-y,x)
------
Rule:
If you start with y = x
2 to the left gives y= x+2
---
then 3 up gives y = x+2+3
or y = x+5
---
then the rotation gives:
x = -y+5
---
Final result is y = -x+5
===============================
Cheers,
Stan H.
==============